OpenStudy (jennyrlz):

How do I turn a number into a degree? i am in pre-calculus, we are using the low of cosine to find the angles for a three (SSS) problem. i got the number but i don't get how they turned it into a degree. the answer is: .4815, but they then turn it into a degree of 61.2.... how?

OpenStudy (jhannybean):

Like a radians to degrees conversion?

OpenStudy (jhannybean):

I'll give you an example, say you have \(4\pi\)

OpenStudy (jennyrlz):

Hmm I have to find the angles of the triangle with the given sides of a=8, b= 3 and c=9

OpenStudy (jhannybean):

To turn \(4\pi\) into degrees you multiply \(4\pi \cdot \dfrac{180^\circ}{\pi}\)

OpenStudy (jhannybean):

Ok, can you draw our your triangle?

OpenStudy (jennyrlz):

I alredy drew it, and I have the answer I just can't turn the answer into degrees :/

OpenStudy (jennyrlz):

I used cosA=(b^2+c^2-a^2)/2bc

OpenStudy (jennyrlz):

And got .4815 but how do I turn this into degrees?

OpenStudy (jennyrlz):

The book got 61.2 degreees but I do not know how they got this >.<

OpenStudy (jhannybean):

law of cosines: \(a^2 = b^2 +c^2 -2bc\cos(A) \implies \cos(A) = \dfrac{a^2-(b^2+c^2)}{-2bc}\) where \(a=8~,~ b = 3~,~ c = 9\) Right?

OpenStudy (jhannybean):

Oh, I see, you distributed the negative sign.

OpenStudy (jennyrlz):

My book has cosA=b^2+c^2-a^2 over 2bc

OpenStudy (jhannybean):

Yeah, i didnt completely simplify, I guess that would make it a little different.

OpenStudy (jhannybean):

\[a^2 = b^2 +c^2 -2bc\cos(A) \implies \cos(A) = \dfrac{-a^2+(b^2+c^2)}{2bc}\]\[a=8~,~ b= 3~,~ c = 9\]Plugging these values in I also get \(\cos(A) \approx 0.48148\)

OpenStudy (jhannybean):

Now to get the angle \(A\),, you need tot ake the inverse cosine of the function.

OpenStudy (jhannybean):

that means, \(\cos^{-1}(0.48148) = A\)

OpenStudy (jhannybean):

Setting your calculator in degree mode, and calculating \(\cos^{-1}(\#) \), you should get \(\approx 61.2^\circ\)`

OpenStudy (jennyrlz):

Hmm I got 28.78 degrees

OpenStudy (jennyrlz):

Oh never mind I plugged it in wrong thanks a bunch :D!!!

OpenStudy (jhannybean):

No problem :)

OpenStudy (jennyrlz):

This is my first time using this website :D, I should come more often!

OpenStudy (jhannybean):

yeah! Of course :D This place has really helpful people that help you understand a bunch of questions using various techniques

OpenStudy (jennyrlz):

Thanks again (((o(*゚▽゚*)o)))

OpenStudy (jhannybean):

No problem :)

OpenStudy (jennyrlz):

Does this just stay open or am I supposed to close it 0.o

OpenStudy (jhannybean):

If your question is complete or if you'd like to ask another one, then you should close this one and post up a new question in a new thread :)

OpenStudy (jennyrlz):

Alright, thanks again :p